SUPEROXIDE PHOTOELECTRON ANGULAR DISTRIBUTIONS: VIBRATIONAL DEPENDENCE AS A CONSQUENCE OF BORN-OPPENHEIMER BEHAVIOR

\underline{RICHARD MABBS}, MATTHEW VAN DUZOR, FOSTER MBAIWA and JIE WIE, Department of Chemistry, Washington University, One Brookings Dr., Campus Box 1134 Saint Louis, Missouri 63130, USA; STEPHEN GIBSON, STEVEN CAVANAGH and B. R. LEWIS, Research School of Physical Sciences and Engineering, The Australian National University, Canberra, Australian Capital Territory 0200, Australia.

For atomic anions the photoelectron angular distribution (PAD) is a signature of the bound excess electron wavefunction. However, even for relatively simple molecular anions the relationship between the detachment orbital and the PAD is more complicated. We demonstrate this with a series of near threshold photoelectron images for superoxide detachment, recorded over small photon energy increments. The highest occupied, π_{g} orbital has a strong resemblance to an atomic d-orbital. Detachment from the superoxide anion via the O_{2}(X ^{3}Σ_{g}^{−}, v′)←O_{2}^{−}(X ^{2}Π_{g}, v′′ = 0) transitions reveals that the electron kinetic energy dependence of the PAD resembles that of atomic anions. However, there is also a clear dependence on the vibrational energy deposited in the neutral O_{2}. Comparison with theoretical calculations reveals that this dependence is not due to non-adiabatic effects. Instead it arises in the change of the electronic transition moment across the X−X band, a direct consequence of the parametric dependence of the π_{g} orbital function on the internuclear separation.